# Exploring Linear Attention Alternative for Single Image Super-Resolution*

### v1

@@ Abstract
Deep learning-based single-image super-resolution (SISR) technology focuses on enhancing low-resolution (LR) images into high-resolution (HR) ones. Although significant progress has been made, challenges remain in computational complexity and quality, particularly in remote sensing image processing. To address these issues, we propose our Omni-Scale RWKV Super-Resolution (OmniRWKVSR) model which presents a novel approach that combines the Receptance Weighted Key Value (RWKV) architecture with feature extraction techniques such as Visual RWKV Spatial Mixing (VRSM) and Visual RWKV Channel Mixing (VRCM), aiming to overcome the limitations of existing methods and achieve superior SISR performance. This work has proved able to provide effective solutions for high-quality image reconstruction. Under the 4x Super-Resolution tasks, compared to the MambaIR model, we achieved an average improvement of 0.26% in PSNR and 0.16% in SSIM.
@@ I Introduction
This study proposes a novel approach that combines the RWKV architecture with advanced feature extraction techniques and verified through ablation experiments, aiming to overcome the limitations of existing attention-based methods, such as the quadratic complexity. The main contributions of this study are as follows:
I-1 Feed Forward Network
ChannelMix is superior to traditional MLP in enhancing channel-wise feature mixing and improving information flow, thus outperforming MLP in feature extraction and image reconstruction quality.
I-2 Omni-Quad Shift
@@
I-3 Weighted Key Value (WKV) 2D Scanning
Conventional linear models can only capture the one-dimensional information, which is also quoted ”scanning”. Despite the 2D scanning of Spate-Space model, the quality and performance still demand improving. To address this issue, the 2D scanning version of the WKV mechanism is proposed in this paper.
I-4 Faster Training With SOTA Performance
Our model reaches the highest score on popular datasets, such as Set14, BSD100 and so on with approximately 15% less training time compared to MambaIR.
@@ III-B Visual RWKV Group
A VRG is a collection of residual blocks that integrates the VRSM and VRCM as its core components. The VRSM facilitates the blending of feature maps across different resolutions, while the VRCM confines these features to specific channels. A VRG is composed of multiple Visual RWKV Residual Blocks (VRBs), and the overall structure of the VRB is shown in Fig. 2.
@@ IV-D Ablation study
In the field of SISR, understanding the impact of individual components within a model is crucial for optimizing performance. To this end, we conducted a systematic ablation study on our proposed OmniRWKVSR, focusing on three key components: the feed forward network, the shift mechanism, and the WKV2D scan module.
@@
Quad-Omni Shift 30.1390 0.8458
@@ IV-D2 Shift Mechanism
The shift mechanism is another critical component of our model, and we evaluated three different mechanisms: OmniShift[33], QShift[34], and Quad-Omni Shift. OmniShift captures local spatial dependencies through uniform shifting but has limitations in capturing long-range dependencies. QShift enhances the ability to capture both local and global spatial dependencies by applying different shifts to different parts of the input features. Quad-Omni Shift combines the strengths of both, utilizing a multi-scale shifting strategy that effectively manages various spatial dependencies. In our evaluation using the Set14 dataset, Quad-Omni Shift achieved a PSNR of 30.1390 and an SSIM of 0.8458, outperforming both OmniShift (PSNR 30.1034, SSIM 0.8453) and QShift (PSNR 30.0965, SSIM 0.8411) under test on Set14 dataset, demonstrating its effectiveness in capturing multi-scale features and spatial transformations.
@@ V Conclusion
Although our model shows superior performance in terms of PSNR and SSIM, there are still issues such as texture errors and lack of clarity in the visualized results. Moreover, the performance on the GMH dataset is not satisfactory. Future work will focus on addressing these problems.

### v2

@@ Abstract
Deep learning-based single-image super-resolution (SISR) technology focuses on enhancing low-resolution (LR) images into high-resolution (HR) ones. Although significant progress has been made, challenges remain in computational complexity and quality, particularly in remote sensing image processing. To address these issues, we propose our Omni-Scale RWKV Super-Resolution (OmniRWKVSR) model which presents a novel approach that combines the Receptance Weighted Key Value (RWKV) architecture with feature extraction techniques such as Visual RWKV Spatial Mixing (VRSM) and Visual RWKV Channel Mixing (VRCM), aiming to overcome the limitations of existing methods and achieve superior SISR performance. Our work demonstrates the ability to provide effective solutions for high-quality image reconstruction. Under the 4x Super-Resolution tasks, compared to the MambaIR model, we achieved an average improvement of 0.26% in PSNR and 0.16% in SSIM.
@@ I Introduction
This study proposes a novel approach that combines the RWKV architecture with advanced feature extraction techniques and verified through ablation experiments, aiming to overcome the limitations of existing attention-based methods, such as the quadratic complexity. The main contributions of this study are as follows.
I-A Methodological Innovations
I-A1 Feed Forward Network
ChannelMix is superior to traditional MLP in improving channel-wise feature mixing and improving information flow, thus outperforming MLP in feature extraction and image reconstruction quality.
I-A2 Omni-Quad Shift
@@
I-A3 Weighted Key Value (WKV) 2D Scanning
Conventional linear models can only capture the one-dimensional information, which is also quoted ”scanning”. Despite the 2D scanning of state-space model, the quality and performance still demands improvement. To address this issue, we propose the 2D scanning version of the WKV mechanism in this paper.
I-B Experimental Advances
I-B1 Faster Training With SOTA Performance
Our model reaches the highest score on popular datasets, including Set14, BSD100, and Urban100 with approximately 15% less training time compared to MambaIR.
@@ III-B Visual RWKV Group
A VRG is a collection of residual blocks that integrates the VRSM and VRCM as its core components. The VRSM facilitates the blending of feature maps across different resolutions, while the VRCM confines these features to specific channels. VRCM replaces traditional MLP by computing cross-channel weights via depthwise convolution. A VRG is composed of multiple Visual RWKV Residual Blocks (VRBs), and the overall structure of the VRB is shown in Fig. 2.
@@ IV-D Ablation study
In the field of SISR, understanding the impact of individual components within a model is crucial for optimizing performance. To this end, we conducted a systematic ablation study on our proposed OmniRWKVSR, focusing on three key components: the feed forward network, the shift mechanism, and the WKV2D scan module. And the ablation studies used Set14 (X4).
@@
Omni-Quad Shift 30.1390 0.8458
@@ IV-D2 Shift Mechanism
The shift mechanism is another critical component of our model, and we evaluated three different mechanisms: OmniShift[33], QShift[34], and Omni-Quad Shift. OmniShift captures local spatial dependencies through uniform shifting but has limitations in capturing long-range dependencies. QShift enhances the ability to capture both local and global spatial dependencies by applying different shifts to different parts of the input features. Omni-Quad Shift combines the strengths of both, utilizing a multi-scale shifting strategy that effectively manages various spatial dependencies. In our evaluation using the Set14 dataset, Omni-Quad Shift achieved a PSNR of 30.1390 and an SSIM of 0.8458, outperforming both OmniShift (PSNR 30.1034, SSIM 0.8453) and QShift (PSNR 30.0965, SSIM 0.8411) under test on Set14 dataset, demonstrating its effectiveness in capturing multi-scale features and spatial transformations.
@@ V Conclusion
Although our model shows superior performance in terms of PSNR and SSIM, there are still issues such as texture errors and lack of clarity in the visualized results. Moreover, the performance on the GMH dataset is not satisfactory. OmniRWKVSR underperforms on GMH due to its unique textures (e.g., grassland holes). Future work will integrate texture-adaptive losses.